OpenStudy (annie12m):

Medal if you solve correctly! 4(x – 3) = 7x + 24

OpenStudy (mathmale):

Annie, perhaps you didn't know that it is against the rules to ask others to do your work for you and to hand you the answer. If you want to learn how to do this kind of problem yourself, you've got to become involved in one or more ways. Review the code of Conduct: http://openstudy.com/code-of-conduct. 4(x – 3) = 7x + 24 is to be solved for x. We need to find a way to get all of the x terms on the left and all of the constant terms on the right. How would y ou go about doing that? Hint: start by doing the multiplication in the left side of your equation.

OpenStudy (annie12m):

4(x – 3) = 7x + 24 4x - 12 = 7x +24?